A manufacture of personal computers has an inventory of 10,000 back-up storage drives that sold for $100 per unit last year. The current market Price of these drives is now $70 per unit. By adding one of these drives to their stock of personal computer, the price of each computer is increase by $80 per unit. Should the driver be added ? what is the opportunity cost of these drivers? Explain
1
Expert's answer
2018-03-07T09:42:07-0500
First, the costs associated with a lost or missed opportunity refer to opportunity cost.
Inventory 10,000 * 100 = 1,000,000 for last year.
Currently; 10,000* 70 = 700,000
If they do not add the drives: they will currently loose; 1,000,000 less 700,000 which is $300,000.
If they add the drives to their PCs:
10,000* 150= $1,500,000
They will gain; 1,500,000 minus 1,000,000 which is $500,000 profit.
Therefore, they should add the drivers.